OpenStudy (yamyam70):

FInd the solution set for..help! sin^2 x = sin x

OpenStudy (yamyam70):

not for help.

OpenStudy (anonymous):

sin^2x - sinx = 0 sinx ( sinx - 1 )= 0 hence , sinx = 0 , and sinx = 1 x = 0 , pi, pi/2

OpenStudy (anonymous):

sin^2x - sin x = 0 => sin x( sin x - 1) = 0 => sin x = 0 or sin x-1 = 0 => sin x = 1. Can you figure it out from there?

OpenStudy (yamyam70):

oh we factor it!

OpenStudy (yamyam70):

thanks @pgpilot326 @tanjeetsarkar96 :))

OpenStudy (anonymous):

yep

OpenStudy (yamyam70):

inverse sine of 1 is 90. , why pi / 2 ?

OpenStudy (phi):

90º = pi/2 radians

OpenStudy (yamyam70):

oh yeah, where did he get pi then ? x = 0 , pi, pi/2

OpenStudy (anonymous):

thats in radians .

OpenStudy (phi):